# Phase transitions and spontaneous symmetry breaking in the renormalized Ginzburg-Landau theory

### v1

@@ Abstract
Phase transitions; symmetry breaking; low-dimensional structures; Ginzburg-Landau theory.
@@ I Introduction
The miniaturization of devices has led to several discoveries and the creation of new functions resulting from the microscopic characteristics of materials. For example, decreasing the thickness of the bulk material allows for the downsizing of other dimensions, thereby enabling a higher density of materials in a single direction or plane. The interest of this work lies in understanding and designing materials with various geometries and symmetries from the perspective of phase transitions.
Symmetry breaking in physics refers to a phenomenon where a symmetric and disordered state collapses into an ordered but less symmetric state. In certain cases, this transition represents a bifurcation that a particle can undergo as it moves toward a lower energy state, as illustrated in Fig. 1. When considering non-abelian aspects, symmetry breaking results in degeneracy in the spectrum, since the states in the theory furnish a representation of the symmetry. Spontaneous symmetry breaking is a key concept in statistical field theory and plays a crucial role in Landau’s theory of phase transitions and critical phenomena [1, 2, 3].
Figure 1: Spontaneous symmetry breaking. a) Unbroken symmetry, b) proposed ${\mathbb{Z}_{2}}$-Symmetric system with two possible states, c) Broken symmetry.
Symmetry arguments are powerful tools for deducing critical qualitative properties of phase transitions. The Landau model assumes that a thermodynamic potential $H_{L}[\phi]$ can be expressed as a series in terms of the order parameter $\phi$. This model analyzes the symmetry properties of different phases and has been extensively studied, as referenced in the review by P. C. Hohenberg and A. P. Krekhov [4]. In this paper, we will focus on transitions characterized by order parameters where the symmetry group of the least symmetric phase (the ordered phase) is a subgroup of the symmetry group of the most symmetric phase (the disordered phase). Specifically, we will examine cases where the order parameter is continuous, leading to second-order transitions. In these situations, the free energy $H_{L}[\phi]$ is analytic in terms of $\phi$ and temperature for a spatially uniform system. Near the transition temperature, this free energy can be expressed in a specific form,
@@
$\bullet$ ${\phi=0}$ represents a relative maximum, leading to two other stable equilibrium positions, ${\phi}_{\pm}=\pm{{\phi_{0}}}$, as illustrated in Figs. 1(b) and 1(c). When $r<0$, the $\phi^{2}$ term in Eq. (1) has a negative quadratic coefficient, making $\phi=0$ a local maximum instead of a minimum. Below the transition temperature $T_{c}$, but close to it, the potential reaches an absolute minimum near $\phi=0$. In this ordered (low-temperature) phase, the system can adopt either of the two degenerate values $\phi_{\pm}=\pm|\phi_{0}|$, which necessitates $r<0$ and $b>0$, as shown in Fig. 2 for temperatures $T<T_{c}$. In fact, $r>0$ results in a discontinuous transition, where $\phi$ shifts from zero to a finite value at $T_{c}$. The same consequence follows for $b<0$. Since $r$ changes sign at $T_{c}$ and regularity has been assumed for all thermodynamic variables, Landau’s basic construction leads to the expression for the quadratic coefficient: $r=r_{0}(T-T_{c})$, where $r_{0}$ is a constant [1, 4].
Figure 2: Order parameter variations with temperature for a second-order transition occur at the transition temperature $T_{c}$, illustrating symmetry breaking behavior before and after $T_{c}$.
In summary, the minimization of the Landau functional without taking into account an external field leads to
$\phi_{0}^{2}=\left\{\begin{array}[]{ll}\vspace{0,5cm}0,&\textrm{for}\phantom{.}\tau>0\phantom{...}\textrm{characterizing the disordered phase}\\ -r/b,&\textrm{for}\phantom{.}\tau<0\phantom{...}\textrm{characterizing the ordered phase}.\\ \end{array}\right.$ (5)
Here, $\tau=\frac{T-T_{c}}{T_{c}}$ is the reduced temperature.
The specific heat $C_{p}$ is expressed by substituting the equilibrium value of the order parameter into the free energy given in Eq. (1). This expression can be factorized as follows:
$C_{p}=\left\{\begin{array}[]{ll}\vspace{0,5cm}C_{0},&\textrm{for}\phantom{...}\tau>0\\ C_{0}+\frac{r_{0}^{2}T}{4b},&\textrm{for}\phantom{...}\tau<0\\ \end{array}\right.$ (8)
In this context, $C_{0}$ represents a constant, while $\Delta C_{0}=\frac{r_{0}^{2}T}{4b}$ indicates the mean-field jump in the specific heat as the temperature falls below the transition temperature $T_{c}$. This specific heat jump varies in prominence depending on the intrinsic properties of certain materials, and it may even vanish entirely in some complex superconducting systems. This review will focus on the phenomenon of the disappearance of the specific heat jump, with the aim of identifying its underlying causes.
@@
$\chi^{-1}=\frac{\partial{h}}{\partial\phi}=r+3b\phi^{2}\phantom{..}\Rightarrow\phantom{.}\chi^{-1}\Big{|}_{h\rightarrow 0}=\left\{\begin{array}[]{ll}\vspace{0,5cm}r,&\textrm{for}\phantom{.}\tau>0\\ -2r,&\textrm{for}\phantom{.}\tau<0.\\ \end{array}\right.$ (11)
@@
There are significant size effects in low-dimensional systems that are negligible in bulk materials. In this review, we develop and summarize ideas that lead to a theoretical model of low-dimensional systems. The outline of the review is organized as follows: In Section II, we present the second-order Ginzburg-Landau theory, highlighting its limitations and the necessity of making corrections through a renormalization process. This process emphasizes the critical role of dimensionality in the manifestation and analysis of transitions. Our theoretical results, computed using typical parameters, demonstrate the importance of considering both dimensionality and thermal fluctuations, confirming that size effects significantly influence the thermodynamic properties of materials. Unlike the standard Landau model, this review aims to provide a corrective term for the quadratic coefficient, incorporating contributions from symmetries and thermal fluctuations. We derive an appropriate expression for this quantity in terms of the system’s correlation functions and scattering state contributions. The spatial variations of the order parameter in a fluctuating system are described by a nonlinear polynomial equation, which takes into account the system’s anisotropy and the dependence of the quadratic coefficient on the intrinsic characteristics of the system. By averaging over the system’s fluctuations, we obtain a macroscopic equation for spatial variations that features a dimension-dependent quadratic coefficient. We derive alternative exact expressions for the renormalized quadratic coefficients that contain more information. In Section III, we highlight evidence that the obtained model allows for better simulation of certain physical conditions involved in phase transitions and predicts specific physical constraints. This is achieved by re-expressing key thermodynamic parameters, notably the jump in specific heat and susceptibility, both of which depend on the system’s dimensionality. Finally, in Section III, we provide a conclusion.
@@ II The renormalized Ginzburg-Landau theory
The basic reason for the failure of Landau’s theory is the neglect of spatial fluctuations, which stems from the assumption that the order parameter is spatially homogeneous. To account for the effects of these spatial fluctuations, the order parameter should be treated as a tensor field instead of just a constant. Consequently, the Landau potential is replaced by an integral over a potential density, known as the Ginzburg-Landau functional [5],
@@
The coefficients $\mathcal{C}$ and $b$ are assumed to be weakly dependent on external thermodynamic parameters (e.g. temperature and pressure).
Although Ginzburg-Landau theory accounts for fluctuations, it has limitations when it comes to describing systems with reduced dimensionality. For example, it cannot adequately explain the order of phase transitions in certain layered bulk materials [6]. Additionally, the theory falls short in addressing one-dimensional [7] and quasi-two-dimensional systems, where new states with unexpected properties emerge at non-zero temperatures [8]. Another significant drawback of Ginzburg-Landau theory is its inability to account for the absence or disappearance of the specific heat jump observed in some materials, such as $YBa_{2}Cu_{3}O_{6+\delta}$ [9], $Ba_{0.2}K_{0.8}Fe_{2}As_{2}$ [10] and certain families of high-$T_{c}$ superconductors, such as $TlBaCaCuO$ and $BiSrCaCuO$ [11].
@@ II.1 Propositions
The importance of critical fluctuation effects, their dynamics, and their impact on the physical properties of materials has been explored in various studies [12, 13, 14, 15, 16, 17, 18, 19]. However, there are still common challenges when it comes to assigning the parameters of the standard Ginzburg-Landau theory to their characteristic quantities, such as temperatures, in a universal way [20]. The renormalization process refines the standard Ginzburg-Landau theory by adding a correction term to the quadratic coefficient, which isn’t accounted for in the standard model. By following this approach, fluctuation effects are incorporated by replacing the quadratic coefficient with its dimensional form
@@
By employing this transformation, we can frame our discussion by qualitatively comparing the results obtained for various dimension values. The model derived from the transformation described in Eq. (7) is phenomenological, as we propose the existence of fluctuations without detailing their origins. The auto-coherent quantity $\Omega(d,T)$ given by [21]:
$\Omega_{d}=\eta\left(\frac{T}{T_{c}}\right)\left(\tau+\frac{\Omega_{d}}{r_{0}T_{c}}\right)^{\frac{d}{2}-1}\phantom{...}\textrm{where}\phantom{...}\tau=\frac{T-T_{c}}{T_{c}},$ (14)
@@
contains important information about regulating the anisotropy of the order parameter. By competing with $r_{0}(T-T_{c})$, the function $\Omega(d,T)$ helps eliminate divergences in field theory [22]. We conceptualize the renormalization process as a collection of techniques designed to address the infinities that arise in calculated thermodynamic quantities. This is achieved by modifying expressions of these quantities to account for the effects of their self-interactions. Using a self-consistent approximation for the quadratic coefficient, we aim to achieve modifications in thermodynamic parameters that incorporate fluctuations, thereby removing the unphysical maximum observed in the crossover regime of standard theory. The inclusion of the $\eta$ term accounts for contributions from scattering states and repulsive interactions between quasi-particles. This competition between $r_{0}(T-T_{c})$ and $\Omega(d,T)$ suggests the coexistence of competing phases within the system, with each phase corresponding to a spontaneously broken symmetry. As a result, the repulsive fluctuation effects cause unsynchronized behavior among quasi-particles, significantly altering the underlying physics [23]. This concept of competing states helps explain the emergence of new states that remain unexplained, depending on whether attractive or repulsive interactions dominate. Fluctuation effects, similar to the orbital effect, have a greater impact on lower phase destruction compared to higher phases. Considering the contributions from scattering states at the Gaussian level clarifies why, in standard Ginzburg-Landau theory, both the dimensional and temperature dependencies of the quadratic coefficient must be taken into account. Even when the paramagnetic effect becomes predominant, this approximation holds true. We will further discuss the significance and role of the energy parameter $\eta$ later.
II.2 Derivation of new quadratic coefficients
@@
To calculate thermodynamic quantities directly from the renormalized Ginzburg-Landau functional, we need to identify the true critical temperature, incorporating the correction term. This critical temperature will be denoted as $T_{c_{d}}$, where the index $d$ indicates the dimension of the system. Using the expression for the correction term from Eq. (8), we can derive a self-consistent equation for the renormalized quadratic coefficient using the following Keumo-Landau polynomial equation [21]:
@@
Upon examining the important result presented in Eq. (9) more closely, we observe that fluctuations increase with $r(d,T)$ raised to the power of $d/2-1$, as long as fluctuation corrections cannot be ignored for dimensions $d\leq 2$. $d_{c}=2$ represents the critical dimension for this approach, below which fluctuations become highly significant. This observation aligns with the Hohenberg-Mermin-Wagner theorem, which states that continuous symmetries cannot be spontaneously broken at finite temperatures in systems with sufficiently short-range interactions when the dimension is $d\leq 2$ [24, 25].
Solutions of Equation (9) as a function of the dimension are given by the following Keumo-Landau quadratic coefficients,
$r(d,T)=\left\{\begin{array}[]{llll}\vspace{0,5cm}r(d=1,T),&\textrm{for}\phantom{.}d=1\\ \eta\left(\frac{T}{T_{c}}\right)+r_{0}(T-T_{c}),&\textrm{for}\phantom{.}d=2,\\ \frac{\eta{T}\Big{(}\eta{T}\pm\sqrt{\eta^{2}T^{2}+4r^{2}_{0}T^{3}_{c}(T-T_{c})}\Big{)}}{2r_{0}T^{3}_{c}}+r_{0}(T-T_{c}),&\textrm{for}\phantom{.}d=3,\\ r_{0}\big{(}1-\frac{\eta{T}}{r_{0}T^{2}_{c}}\big{)}^{-1}(T-T_{c}),&\textrm{for}\phantom{.}d=4.\\ \end{array}\right.$ (20)
@@
Figure 3: The plot illustrates the standard and renormalized quadratic coefficients for a one-dimensional system. It is assumed that $T_{c}=5K$ and $\eta$ = 0.8.. The solid curve represents the correction term $\Omega_{1D}(T)$, the dotted curve shows the standard quadratic coefficient $r_{0}(T-T_{c})$, and the dashed curve depicts the behavior of the renormalized quadratic coefficient, defined as $r(d,T)=r_{0}(T-T_{c})+\Omega_{1D}(T)$. While the standard coefficient goes to zero at $T_{c}$, the renormalized quadratic coefficient asymptotically approaches zero, i.e. $T_{c_{1D}}=0$ is the critical temperature for one-dimensional systems in the renormalized theory
@@
The Keumo-Landau quadratic coefficients still cancel out at a finite transition temperature given by $r(d,T)$ = 0 (except for one-dimensional systems), but the critical behavior is qualitatively and quantitatively modified. However, the transition temperature is no longer monotonic because it depends on the system’s dimensionality.
$r(d,T)=0\Rightarrow\left\{\begin{array}[]{llll}\vspace{0,3cm}T_{c_{1D}}=0,&\textrm{for}\phantom{.}d=1\\ \vspace{0,3cm}T_{c_{2D}}=T_{c}(1+\frac{\eta}{r_{0}T_{c}})^{-1},&\textrm{for}\phantom{.}d=2,\\ \vspace{0,3cm}T_{c_{3D}}=T_{c},&\textrm{for}\phantom{.}d=3,\\ \vspace{0,3cm}T_{c_{4D}}=T_{c},&\textrm{for}\phantom{.}d=4,\\ \end{array}\right.$ (26)

### v2

# Induced phase transitions and spontaneous symmetry breaking based on the renormalized Ginzburg-Landau theory

@@ Abstract
Phase transitions; spontaneous symmetry breaking; renormalization; Ginzburg-Landau theory; low-dimensional structures.
@@ I Introduction
So far, the miniaturization of devices has led to several discoveries and the creation of new functions resulting from the microscopic properties/characteristics of materials. For instance, a higher density of materials in a single direction or plane is made possible by reducing the thickness of the bulk material, which also permits the downsizing of other dimensions. Accordingly, understanding and designing new materials with different geometries and symmetries, which generally result in new functions and demonstrate many intriguing properties, i.e., tremendous characteristics from the standpoint of phase transitions, continue to considerably draw the attention of many researchers [1, 2].
Symmetry breaking in physics refers to a phenomenon where a symmetric and disordered state collapses into an ordered but less symmetric state [3, 4, 2]. In certain cases, this transition represents a bifurcation that a particle can undergo as it moves toward a lower energy state, as illustrated in Fig. 1. When considering non-abelian aspects, symmetry breaking results in degeneracy in the spectrum, since the states in the theory furnish a representation of the symmetry. Amongst other things, spontaneous symmetry breaking is well known as a key concept in statistical field theory and plays a crucial role in the so-called Ginzburg-Landau theory of phase transitions and critical phenomena [1, 3, 4, 5, 2, 6].
Figure 1: Schematic illustrations of the evolution of the effective on-site potential corresponding to: a) non-breaking/Unbroken symmetry, b) proposed $\mathbb{Z}_{2}$}-symmetric system with two possible states, and c) effective symmetry breaking/ Broken symmetry. This general framework identifies the mechanisms and provides the guideline for the phase transition mechanism.
More importantly, symmetry breaking arguments in general are powerful tools to deduce and control critical qualitative properties of phase transitions, further leading to a striking pattern with lower symmetry, i.e., breaking the original permutation symmetries of unit cells [1, 7] and also to macroscopic buckling states [8, 9]. Indeed, the Landau model assumes that a thermodynamic potential $H_{L}[\phi]$ can be expressed as a series in terms of the order parameter $\phi$, and the model analyzes the symmetry properties of different phases and has been extensively studied, as referenced in the review by Hohenberg and Krekhov [7]. In this paper, we will focus on transitions characterized by order parameters where the symmetry group of the least symmetric phase (the ordered phase) is a subgroup of the symmetry group of the most symmetric phase (the disordered phase). Specifically, we will examine cases where the order parameter is continuous, leading to second-order transitions. In these situations, the free energy $H_{L}[\phi]$ is analytic in terms of $\phi$ and temperature for a spatially uniform system. Near the transition temperature, this free energy can be expressed in a specific form,
@@
$\bullet$ $\phi=0$} represents a relative maximum, leading to two other stable equilibrium positions, ${\phi}_{\pm}=\pm{{\phi_{0}}}$, as illustrated in Figs. 1(b) and 1(c). When $r<0$, the $\phi^{2}$ term in Eq. (1) has a negative quadratic coefficient, making $\phi=0$ a local maximum instead of a minimum. Below the transition temperature $T_{c}$, but close to it, the potential reaches an absolute minimum near $\phi=0$. In this ordered (low-temperature) phase, the system can adopt either of the two degenerate values $\phi_{\pm}=\pm|\phi_{0}|$, which necessitates $r<0$ and $b>0$, as shown in Fig. 2 for temperatures $T<T_{c}$. In fact, $r>0$ results in a discontinuous transition, where $\phi$ shifts from zero to a finite value at $T_{c}$. The same consequence follows for $b<0$. Since $r$ changes sign at $T_{c}$ and regularity has been assumed for all thermodynamic variables, Landau’s basic construction leads to the expression for the quadratic coefficient: $r=r_{0}(T-T_{c})$, where $r_{0}$ is a constant [1, 5, 7].
Figure 2: Schematic geometry representation of the evolution of the effective on-site potential with temperature, illustrating spontaneous symmetry breaking behavior at $T_{c}$ and the structural phase transition mechanism of the multistable systems.
In general, the minimization of the Landau functional without taking into account an external field leads to:
[…]
Here, $\tau=\frac{T-T_{c}}{T_{c}}$ is the reduced temperature. The specific heat $C_{p}$ is expressed by substituting the equilibrium value of the order parameter into the free energy given in Eq. (1). This expression can be factorized as follows:
[…]
In this context, $C_{0}$ represents a specific heat constant, while $\Delta C_{0}=\frac{r_{0}^{2}T}{4b}$ indicates the mean-field jump in the specific heat as the temperature falls below the transition temperature $T_{c}$. This specific heat jump varies in prominence depending on the intrinsic properties of certain materials, and it may even vanish entirely in some complex superconducting systems. This review will focus on the phenomenon of the disappearance of the specific heat jump, with the aim of identifying its underlying causes.
@@
[…]
@@
There are significant size effects in low-dimensional systems that are negligible in bulk materials. Therefore, the interest of this work lies in understanding and designing materials with various geometries and symmetries from the perspective of structural phase transitions, and which induce orderly localized configurations. Moreover, as well known in the phase transition theory considered up to now, the spontaneous symmetry breaking came from the phase transition; so we will focus on systems with spontaneous symmetry breaking, so that Ginzburg-Landau theory will once again turn out to be useful. In this context and to be more specific, we will rely on a consideration of a system undergoing a continuous phase transition from a high-temperature symmetric phase to a low-temperature ordered phase in which some symmetry is broken.
In this review, we develop and summarize ideas that lead to a theoretical model of low-dimensional systems. The outline of the review is organized as follows: In Section II, we present the second-order Ginzburg-Landau theory, highlighting its limitations and the necessity of making corrections through a renormalization process. This process emphasizes the critical role of dimensionality in the manifestation and analysis of transitions. Our theoretical results, computed using typical parameters, demonstrate the importance of considering both dimensionality and thermal fluctuations, confirming that size effects significantly influence the thermodynamic properties of materials. Unlike the standard Landau model, this review aims to provide a corrective term for the quadratic coefficient, incorporating contributions from symmetries and thermal fluctuations. We derive an appropriate expression for this quantity in terms of the system’s correlation functions and scattering state contributions. The spatial variations of the order parameter in a fluctuating system are described by a nonlinear polynomial equation, which takes into account the system’s anisotropy and the dependence of the quadratic coefficient on the intrinsic characteristics of the system. By averaging over the system’s fluctuations, we obtain a macroscopic equation for spatial variations that features a dimension-dependent quadratic coefficient. We derive alternative exact expressions for the renormalized quadratic coefficients that contain more information. In Section III, we highlight evidence that the obtained model allows for better simulation of certain physical conditions involved in phase transitions and predicts specific physical constraints. This is achieved by re-expressing key thermodynamic parameters, notably the jump in specific heat and susceptibility, both of which depend on the system’s dimensionality. Accordingly, Section IV presents, we provide a comprehensive discussion of the results and their implications. Finally, Section V offers a conclusion.
@@ II The renormalized Ginzburg-Landau theory
The basic reason for the failure of Landau’s theory is the neglect of spatial fluctuations, which stems from the assumption that the order parameter is spatially homogeneous. To account for the effects of these spatial fluctuations, the order parameter should be treated as a tensor field instead of just a constant. Consequently, the Landau potential is replaced by an integral over a potential free energy density, known as the Ginzburg-Landau Hamiltonian functional [4],
@@
Here, the coefficients $\mathcal{C}$ and $b$ are assumed to be weakly dependent on external thermodynamic parameters (e.g. temperature and pressure).
Although Ginzburg-Landau theory accounts for fluctuations, it has limitations when it comes to describing systems with reduced dimensionality. For example, it cannot adequately explain the order of phase transitions in certain layered bulk materials [10]. Additionally, the theory falls short in addressing one-dimensional [11] and quasi-two-dimensional systems, where new states with unexpected properties emerge at non-zero temperatures [12]. Another significant drawback of Ginzburg-Landau theory is its inability to account for the absence or disappearance of the specific heat jump observed in some materials, such as YBa2Cu3O6+δ [13], Ba0.2K0.8Fe2As2 [14] and certain families of high-$T_{c}$ superconductors, such as TlBaCaCuO and BiSrCaCuO [15].
@@ II.1 Propositions
The importance of critical fluctuation effects, their dynamics, and their impact on the physical properties of materials has been explored in various studies [16, 17, 18, 19, 20, 21, 22, 23]. However, there are still common challenges when it comes to assigning the parameters of the standard Ginzburg-Landau theory to their characteristic quantities, such as temperatures, in a universal way [24]. The renormalization process refines the standard Ginzburg-Landau theory by adding a correction term to the quadratic coefficient, which isn’t accounted for in the standard model. By following this approach, effects of fluctuations and correlations of order parameter are incorporated by replacing the quadratic coefficient with its dimension-dependent expression
@@
By employing this transformation, we can frame our discussion by qualitatively comparing the results obtained for various dimension values. The model derived from the transformation described in Eq. (7) is phenomenological, as we propose the existence of fluctuations without detailing their origins. In this context, the auto-coherent dimension-dependent correction quantity $\Omega(d,T)$, first derived within the framework of the renormalized Gaussian approach by Keumo et al. and which satisfy the following self-consistent equation (see appendix A) [25]:
$\Omega(d,T)=\eta\left(\frac{T}{T_{c}}\right)\left[\tau+\frac{\Omega(d,T)}{r_{0}T_{c}}\right]^{\frac{d}{2}-1},\phantom{...}\textrm{where}\phantom{...}\tau=\frac{T-T_{c}}{T_{c}},$ (14)
contains important information about regulating the anisotropy of the order parameter. This expression shall for historical reasons referred to as ’Keumo-Mkam-Danga-Domngang-Hounkonnou’(KMDDH) correction coefficient relation of the quadratic expansion of the Ginzburg-Landau free energy functional induced by the quantum fluctuations of the order parameter.
Here the coefficient $\eta$ denotes a variational energetic parameter related to the harmonic variance of the order parameter which takes into account the proper dependence of the magnitude and the temperature of the fluctuation corrections on the anisotropy of the electron spectrum and the possibly related anomalies including the impurity concentration, stacking faults and other intrinsic defects shown by physical properties [26, 22, 27]. More specifically, the quantity $\eta$ generally depends on both the coherence length and the fourth order coupling constant which is known to account for the self-interaction of the system and the control of the saturation of ordering in the Ginzburg-Landau theory, and its magnitude can induce or prevent phase transitions [27]. Furthermore, the Ginzburg-Landau Hamiltonian $H_{GL}$, as a function of a spatially varying generalized single order parameter configuration $\phi(\textbf{r})$ is now renormalized, which finally leads us to a coarse-grained effective Ginzburg-Landau Hamiltonian involving fluctuations and in the presence of an external field h, and which may now cast the form
$H_{RGL}[\phi]=\int d^{d}\textbf{r}\bigg[\frac{\mathcal{C}}{2}\big[\nabla_{r}\phi(\textbf{r})\big]^{2}+\frac{r(d,T)}{2}\phi^{2}+\frac{b}{4}\phi^{4}-h\phi\bigg].$ (15)
where the phenomenological renormalized quadratic coefficient $r(d,T)=r_{0}(T-T_{c})+\Omega(d,T)$. This new coefficient takes into account the fluctuations on the two-point correlation function, which implies the destruction of the long-range order, as predicted by the mean-field theory [1, 4, 5, 7].
@@
By competing with $r_{0}(T-T_{c})$ in the renormalized Ginzburg-Landau Hamiltonian Eq. (15), the dimension-dependent function $\Omega(d,T)$ helps eliminate divergences in field theory [28]. We conceptualize the renormalization process as a collection of techniques designed to address the infinities that arise in calculated thermodynamic quantities. This is achieved by modifying expressions of these quantities to account for the effects of their self-interactions. Using a self-consistent approximation for the quadratic coefficient, we aim to achieve modifications in thermodynamic parameters that incorporate fluctuations, thereby removing the unphysical maximum observed in the crossover regime of standard theory. The inclusion of the $\eta$ term accounts for contributions from scattering states and repulsive interactions between quasi-particles. This competition between $r_{0}(T-T_{c})$ and $\Omega(d,T)$ suggests the coexistence of competing phases within the system, with each phase corresponding to a spontaneously broken symmetry. As a result, the repulsive fluctuation effects cause non-synchronized behavior among quasi-particles, significantly altering the underlying physics [27]. This concept of competing states helps explain the emergence of new states that remain unexplained, depending on whether attractive or repulsive interactions dominate. Fluctuation effects, similar to the orbital effect, have a greater impact on lower phase destruction compared to higher phases. Considering the contributions from scattering states at the Gaussian level clarifies why, in standard Ginzburg-Landau theory, both the dimensional and temperature dependencies of the quadratic coefficient must be taken into account. Even when the paramagnetic effect becomes predominant, this approximation holds true. We will further discuss the significance and role of the energy parameter $\eta$ later.
II.2 Dimension-dependent renormalized quadratic coefficient and treatment
@@
In order to illuminate the physics of thermodynamic phases and leads to powerful theoretical methods for understanding critical phenomena at continuous phase transitions, it is important to investigate quite generally the dependence of the critical point on the quantum fluctuations throughout the scaling and fluctuation corrections. In this context, the primary objective of the present study is to investigate the effects of spatial variations and fluctuations on the Ginzburg-Landau functional expansion. Accordingly, to calculate thermodynamic quantities directly from the renormalized Ginzburg-Landau functional, we need to identify the true critical temperature, incorporating the correction term. This critical temperature will be denoted as $T_{c_{d}}$, where the index $d$ indicates the dimension of the system. By using the expression for the KMDDH correction term $\Omega(d,T)$ from Eq. (14), we can derive an equation for the dimension-dependent renormalized quadratic coefficient $r(d,T)$ given by the following self-consistent dimension-dependent KMDDH polynomial equation [25]:
@@
Upon examining the important result presented in Eq. (16) more closely, we observe that fluctuations increase with $r(d,T)$ raised to the power of d/21d/2-1, as long as fluctuation corrections cannot be ignored for dimensions $d\leq 2$. $d_{c}$ = 2 represents the critical dimension for this approach, below which fluctuations become highly significant. This observation aligns with the Hohenberg-Mermin-Wagner theorem, which states that continuous symmetries cannot be spontaneously broken at finite temperatures in systems with sufficiently short-range interactions when the dimension is $d\leq 2$ [41, 30].
By considering the different physical dimension $d$, solutions of Eq. (16) is dimension-dependents and are expressed by the following renormalized quadratic coefficients,
[…]
@@
Figure 3: The plot illustrates the standard and renormalized quadratic coefficients for a one-dimensional system. It is assumed that $T_{c}=5K$ and $\eta$ = 0.8. The solid curve represents the correction term $\Omega_{1D}(T)$, the dotted curve shows the standard quadratic coefficient $r_{0}(T-T_{c})$, and the dashed curve depicts the behavior of the renormalized quadratic coefficient, defined as $r(d,T)=r_{0}(T-T_{c})+\Omega_{1D}(T)$. While the standard coefficient goes to zero at $T_{c}$, the renormalized quadratic coefficient asymptotically approaches zero, i.e. $T_{c_{1D}}$ = 0 is the critical temperature for one-dimensional systems in the renormalized theory.
@@
It is important to precise that, the obtained renormalized quadratic coefficients still cancel out at a finite transition temperature given by $r(d,T)$ = 0 (except for one-dimensional systems), but the critical behavior is qualitatively and quantitatively modified. However, the transition temperature is no longer monotonic because it depends on the system’s dimensionality.
[…]